# DiFFPO: Training Diffusion LLMs to Reason Fast and Furious via Reinforcement Learning

### v1

@@ 5 Related Works
RL for dLLMs. Comparing to the extensive study on RLHF (Ouyang et al., 2022; Winata et al., 2025) and RLVR (Lambert et al., 2024; Zhang et al., 2025) for LLMs, there is noticably less study on dLLMs. d1 (Zhao et al., 2025) proposes diffu-GRPO, which is the first work on training base dLLMs for reasoning with RL. Earlier work like Zekri & Boullé (2025) has explored policy gradient methods on smaller scale models. Zhu et al. (2025) proposed a DPO-style (Rafailov et al., 2023) preference optimization method underpinning the development of LLaDA 1.5. Yang et al. (2025) proposed UniGRPO which utilizes random partial masking on completions instead of full masking. DiffuCoder (Gong et al., 2025) used coupled-GRPO to improve the base coding dLLMs. LLaDOU (Huang et al., 2025) trained a token-position aware sampler for enhanced dLLMs reasoning performance. Concurrent to our work, TraceRL (Wang et al., 2025c) also proposes to utilize online trajectory token latents, which is similar to our two times mean-field approximation, yet our parameterization based on time steps is more generalizable and flexible for different dLLM samplers.

### v2

@@ 5 Related Works
RL for dLLMs. Comparing to the extensive study on RLHF (Ouyang et al., 2022; Winata et al., 2025) and RLVR (Lambert et al., 2024; Zhang et al., 2025) for LLMs, there is noticably less study on dLLMs. Zekri and Boullé (2025) proposed to fine tune smaller scale discrete diffusion models with policy gradient methods, yet it requires full likelihood computation, which is hard and expensive for modern dLLMs. d1 (Zhao et al., 2025) proposes diffu-GRPO, which is the first work on training base dLLMs for reasoning with RL. Zhu et al. (2025) proposed a DPO-style (Rafailov et al., 2023) preference optimization method underpinning the development of LLaDA 1.5. Yang et al. (2025) proposed UniGRPO which utilizes random partial masking on completions instead of full masking. LLaDOU (Huang et al., 2025) trained a token-position aware sampler for enhanced dLLMs reasoning performance. Concurrent to our work, wd1 (Tang et al., 2025) used weighted advantage function to improve upon $d1$ on the same math and planning tasks as ours, DiffuCoder (Gong et al., 2025) used coupled-GRPO to improve the base coding dLLMs, TraceRL (Wang et al., 2025c) also proposes to utilize online trajectory token latents, which is similar to our two times mean-field approximation, yet our parameterization based on time steps is more generalizable and flexible for different dLLM samplers.
@@ 6 Discussion and Future Works
Appendix A More experimental results
A.1 Benefits of Two Times Mean Field Approximation
We empirically measure the evolution of average KL divergence during the iterative decoding process of dLLMs across multiple reasoning benchmarks and the evidence that our proposed two times mean field approximation provides better approximation.
Experimental Setup. We use the LLaDA instruct version for inference and use confidence based remasking. We set the number of allowed unmasked token per time step to be 1. Temperature is set as 0 for deterministic generation. For each dataset, we choose 100 prompts to compute the average and set block size to 32. For simplicity, we only compute the results for the first block.
Results on different datasets. As showcased in Figure, for all datasets (GSM8K, Countdown, Sudoku and MATH), we have reduced avg KL divergence error by using two times mean field approximation.
Figure 6: KL divergence between approximated likelihood and true likelihood.
A.2 Importance ratio clamping hyperparameter sweeping
Importance ratio clamping. We also study the effect of the role of importance ratio clamp hyperparameter $C$ and compute the average value after clamping for different clamping values. The results are showcased in Figure 7. Overall the smaller the clamp constant, the smaller the average of the clamped importance ratio. Despite that the constants in our clamp value search space is all able to control extreme ratio values, we still witness that the constants will have an non neglectable effect on the final obtained algorithm/model performance.
Figure 7: KL divergence between approximated likelihood and true likelihood.